# JointDiff: Bridging Continuous and Discrete in Multi-Agent Trajectory Generation

### v1

@@ 4.3.1 CrossGuid for Controllable Generation
The CrossGuid module is integrated within each Social-Temporal Block, situated between the Temporal Mamba and the first Social Transformer. It refines the intermediate representation $H\in\mathbb{R}^{T\times N\times 256}$ (obtained after processing by the Temporal Mamba and Layer Normalization) using the conditioning tensor $G$. The operation is defined as $H^{\prime}=\textrm{CrossGuid}(H,G)=H+\Delta H$, where the residual update $\Delta H$ is computed via a Multi-Head Attention (MHA) mechanism (Vaswani et al., 2017). The specific implementation varies with the guidance modality, as detailed below and shown in Fig. 2-right, where $H^{\top},H^{\prime\top}\in\mathbb{R}^{N\times T\times 256}$ are the transposed tensor of $H$ and $H^{\prime}$, respectively.
@@
Text-Guidance. This modality conditions generation on a natural language prompt ${\mathcal{G}}_{\text{text}}$. The text is tokenized and encoded using a frozen, pre-trained T5-Base encoder (Raffel et al., 2020), producing $G\in\mathbb{R}^{L\times 768}$.
@@ 5 Experiments
Continuous trajectories ($\bm{\mathrm{Y}}$). We validate JointDiff on three public sports datasets: NBA, NFL, and BundesLiga. The NBA dataset uses the widely adopted SportVU data [^1] https://github.com/linouk23/NBA-Player-Movements, with the splits from Mao et al. (2023) (32.5k training / 12k testing scenes). Each scene spans 6 seconds ($T=30$ timesteps, 5 fps) with $N=11$ agents (the ball and 10 players). The NFL dataset comes from the Big Data Bowl [^2] https://github.com/nfl-football-ops/Big-Data-Bowl, following the splits of Xu & Fu (2025) (10,762 training / 2,624 testing scenes). Each scene covers 5 seconds ($T=50$, 10 fps) with $N=23$ agents (the ball and 22 players). Finally, the BundesLiga dataset is curated from the German soccer league [^3] https://github.com/spoho-datascience/idsse-data (Bassek et al., 2025), containing 2,093 training and 524 testing scenes from 7 matches. Scenes with fewer than $N=23$ agents or out-of-play were removed. Each scene spans 6.4 seconds ($T=40$ timesteps, 6.25 fps) with $N=23$ agents, and the training set is augmented with 180 rotations, doubling its size.
@@
Implementation. We use $S=50$ diffusion steps for continuous and $S^{d}=10$ for discrete data, with a shared quadratic noise scheduler from $\beta_{0}=10^{-4}$ to $\beta_{S}=0.5$ (Capellera et al., 2025; Tashiro et al., 2021). The discrete loss coefficient is set to $\lambda=0.1$, following an ablation study in Appendix D.3.2. Sampling employs DDIM with a skip interval $\zeta=5$ and an extra denoising step at $s=1$, yielding 11 steps: $\{50,45,\dots,5,1,0\}$. Training runs for 100 epochs (NBA/NFL) and 200 epochs (BundesLiga) with batch size 16; learning rate $10^{-3}$ is halved every 20 (NBA/NFL) or 40 (BundesLiga) epochs. The model uses a hidden size of 256 and 8 attention heads in all multi-head attention layers, while the Social Transformer employs a 1024-dimensional feedforward layer.
@@ 5.1 Completion Generation
In the imputation generation setting, models are provided with the first 10 frames and the final frame, and must predict the missing in-between trajectories. We evaluate against the IID models Sports-Traj (Xu & Fu, 2025), U2Diff, and the deterministic uni-modal TranSPORTmer (Capellera et al., 2024), reporting the SADE metric. As shown in Table 1-bottom, our method achieves the SOTA in this benchmark. The results show that diffusion-based models, such as U2Diff (Capellera et al., 2025) and Ours, consistently outperform the CVAE-based model Sports-Traj (Xu & Fu, 2025), which suffers from mode collapse. Additional quantitative results, including human evaluation and qualitative results are provided in Appendix D.2 and  D.5.1, respectively.
@@ 6 Conclusions
Acknowledgment. This work has been supported by the project GRAVATAR PID2023-151184OB-I00 funded by MCIU/AEI/10.13039/501100011033 and by ERDF, UE and by the Government of Catalonia under 2023 DI 00058.
@@ Appendix B Architecture
Figure 4: Model Architecture (extended version).
@@
  • Temporal Mamba (TM): As described in Capellera et al. (2025), TM applies bidirectional Mamba modules (Gu & Dao, 2023) independently to each agent and sums the results.
@@ C.1 Possessor threshold
Our methodology is based on the key observation that a ball’s trajectory is primarily linear when it is not in any player’s possession (e.g., during a pass or a shot). Conversely, a player taking possession or influencing the ball’s path will induce a significant change in its direction.
To quantify this, we analyzed the change in the ball’s direction by computing the angle between consecutive ball velocity vectors across our training datasets. For a range of distance thresholds from 0 to 3 meters, we calculated the average angle of change for the ball only during periods when it was outside of that specific threshold distance from all players. The results are shown in Fig. 5.
@@ C.2.1 NFL Dataset
  • Data Source: We process raw tracking data from the NFL Big Data Bowl, adopting the data splits and experimental settings from Xu & Fu (2025).
@@ D.1 Baselines
  • Sports-Traj (Xu & Fu, 2025): We used the official repository with the same hyperparameters as in their benchmark. We observed, consistent with the authors’ checkpoints in another benchmark, that the 20 sampled modes were nearly identical.
@@ D.2.1 Human-based Metrics
To assess the perceptual quality of our generated scenes, we conducted a human evaluation using a pairwise comparison methodology on the NBA dataset. Fifteen participants were asked to select their preferred scene from 15 randomly paired generations from our model, JointDiff, and three baselines: our variant w/o joint, U2Diff, and MoFlow. The Ground Truth (GT) was also included as a reference point.
To ensure a fair evaluation of each model’s best-case performance, we first generated 20 candidate modes for each model. The mode with the minimum SADE metric was then selected as the representative sample. For each of the first 128 past observations in our test set, we formed a random pair from these selected samples. An example of the interface is shown in Fig. 6.
The winning rate of our model against the baselines is shown in Fig. 7. JointDiff consistently outperforms the baselines, achieving higher win rates against MoFlow (80%), U2Diff (65%), and our ablated version w/o joint (53%). The high win rate against the baselines, coupled with an expected loss rate against the GT, indicates that JointDiff generates high-quality, realistic trajectories that are perceptually superior to those produced by other methods. Note that excluding ties, the win rate against our ablated version is increases to 67%, meaning that our proposed joint approach is affine to the human eye. Also the amount of loss against GT is less than the half 44% and the users seems to be confused in a significant amount of samples, 24%.
@@
Figure 7: Winning Rate. Human evaluation results based on pairwise comparisons in NBA generated data. The histogram shows the proportion of wins, ties, and losses for JointDiff versus each competitor, with $n$ indicating the total number of comparisons for each pair.
@@ D.2.3 Completion generation
We benchmark our approach on general trajectory completion using the datasets and experimental setup from Xu & Fu (2025), which includes the Basketball-U, Football-U, and Soccer-U datasets. In this task, a pre-defined mask using different strategies is applied to a scene, and the model must complete the missing observations. We note that while our NFL dataset uses the same data splits as Football-U, it employs a different masking strategy. For comparison, we reuse the evaluation table from Xu & Fu (2025), also reported in Capellera et al. (2025), which includes a wide range of baselines. These baselines span statistical methods: Mean, Median, Linear Fit; vanilla models: LSTM (Hochreiter & Schmidhuber, 1997), Transformer (Vaswani et al., 2017)); and advanced methods: MAT (Zhan et al., 2019), Naomi (Liu et al., 2019), INAM (Qi et al., 2020), SSSD (Alcaraz & Strodthoff, 2022), GC-VRNN (Xu et al., 2023), Sports-Traj (Xu & Fu, 2024), and U2Diff (Capellera et al., 2025).
Results are reported in Table 4. We use the ADE metric as defined in Xu & Fu (2025), which we rename as BADE to reflect its dependence on batch size and distinguish it from the standard individual-level ADE or the scene-level SADE. The BADE metric is defined as:
@@
where $y^{b}_{t,n}$ is the ground truth 2D spatial location of agent $n$ at timestep $t$ in scene $b$, $\hat{y}^{b}_{t,n}$ is its estimation, and $m_{t,n}^{b}$ is a value from the binary mask $\mathbf{M}$ where a value of 0 indicates a location to be predicted. In their setting, Xu & Fu (2025) use $B=128$. We also report the standard SADE metric. The table presents the minimum (minmin) values across 20 generated modes. While the minimum mode for BADE is selected across the entire batch of scenes, the minimum mode for SADE is selected for each scene independently, making it independent of batch size and mode ordering.
@@
Table 4: Completion Generation. We report the minmin for BADE (ADE in (Xu & Fu, 2025)) and SADE over 20 generated modes.
@@ D.3.2 Lambda in Joint Loss
D.3.3 Additional Analysis
We present the last ablation analysis which includes the relevance of two important components in our approach. The first one is the importance sampling from Nichol & Dhariwal (2021), and the second is the number of Social Transformers layers within residual denoising block. Table 7 shows the results of our approach without the importance sampling (w/o IS) and with only one Social Transformer layer (w 1 $\times$ ST). Notice that the importance sampling is crucial when the size of the dataset is small, like the NFL and the BundesLiga. Using two Social Transformers also improve the results across the three datasets.

### v2

@@ 4.3.1 CrossGuid for Controllable Generation
The CrossGuid module is integrated within each Social-Temporal Block, situated between the Temporal Mamba and the first Social Transformer. It refines the intermediate representation $H\in\mathbb{R}^{T\times N\times 256}$ (obtained after processing by the Temporal Mamba and Layer Normalization) using the conditioning tensor $G$. The operation is defined as $H^{\prime}=\textrm{CrossGuid}(H,G)=H+\Delta H$, where the residual update $\Delta H$ is computed via a Multi-Head Attention (MHA) mechanism (Vaswani et al., 2017). The implementation varies with the guidance modality, as detailed below and shown in Fig. 2-right, where $H^{\top},H^{\prime\top}\in\mathbb{R}^{N\times T\times 256}$ are the transposed tensor of $H$ and $H^{\prime}$, respectively.
@@
Text-Guidance. This modality conditions generation on a natural language prompt ${\mathcal{G}}_{\text{text}}$. The text is tokenized and encoded using a frozen, pre-trained T5-Base Encoder (Raffel et al., 2020), producing $G\in\mathbb{R}^{L\times 768}$, where $L$ is the number of tokens and $d=768$.
@@ 5 Experiments
Continuous trajectories ($\bm{\mathrm{Y}}$). We validate JointDiff on three public sports datasets: NBA, NFL, and Bundesliga. The NBA dataset uses the widely adopted SportVU data [^1] https://github.com/linouk23/NBA-Player-Movements, with the splits from Mao et al. (2023) (32.5k training / 12k testing scenes). Each scene spans 6 seconds ($T=30$ timesteps, 5 fps) with $N=11$ agents (the ball and 10 players). The NFL dataset comes from the Big Data Bowl [^2] https://github.com/nfl-football-ops/Big-Data-Bowl, following the splits of Xu and Fu (2025) (10,762 training / 2,624 testing scenes). Each scene covers 5 seconds ($T=50$, 10 fps) with $N=23$ agents (the ball and 22 players). Finally, the Bundesliga dataset is curated from the German soccer league [^3] https://github.com/spoho-datascience/idsse-data (Bassek et al., 2025), containing 2,093 training and 524 testing scenes from 7 matches. Scenes with fewer than $N=23$ agents or out-of-play were removed. Each scene spans 6.4 seconds ($T=40$ timesteps, 6.25 fps) with $N=23$ agents, and the training set is augmented with 180 rotations, doubling its size.
@@
Implementation. We use $S=50$ diffusion steps for continuous and $S^{d}=10$ for discrete data, with a shared quadratic noise scheduler from $\beta_{0}=10^{-4}$ to $\beta_{S}=0.5$ (Capellera et al., 2025; Tashiro et al., 2021). The discrete loss coefficient is set to $\lambda=0.1$, which provided the best trade-off between trajectory and event accuracy (see Appendix D.3.2). Sampling employs DDIM with a skip interval $\zeta=5$ and an extra denoising step at $s=1$, yielding 11 steps: $\{50,45,\dots,5,1,0\}$. Training runs for 100 epochs (NBA/NFL) and 200 epochs (Bundesliga) with batch size 16; learning rate $10^{-3}$ is halved every 20 (NBA/NFL) or 40 (Bundesliga) epochs. The model uses a hidden size of 256 and 8 attention heads in all multi-head attention layers, while the Social Transformer employs a 1024-dimensional feedforward layer. All models are trained on a single RTX A6000.
@@ 5.1 Completion Generation
In the imputation generation setting, models are provided with the first 10 frames and the final frame, and must predict the missing in-between trajectories. We evaluate against the IID models Sports-Traj (Xu and Fu, 2025), U2Diff, and the deterministic uni-modal TranSPORTmer (Capellera et al., 2024), reporting the SADE metric. As shown in Table 1-bottom, our method achieves the SOTA in this benchmark. The results show that diffusion-based models, such as U2Diff (Capellera et al., 2025) and Ours, consistently outperform the CVAE-based model Sports-Traj (Xu and Fu, 2025), which suffers from mode collapse. Additional quantitative and qualitative results are provided in Appendix D.2 and  D.5.1, respectively.
Figure 3: Human evaluation on NBA future generation. The histogram reports the proportions of wins, ties, and losses for JointDiff against each baseline, with $n$ denoting the number of pairwise comparisons.
To assess perceptual quality, we conduct a human evaluation using pairwise comparisons on the NBA future generation task. Fifteen participants judged fifteen random pairs drawn from JointDiff, our ablated variant without joint modeling (“w/o joint”), U2Diff, MoFlow, and Ground Truth (GT) (interface shown in Appendix D.2.1). For fairness, each model generated 20 modes from the same past, and the sample with the lowest SADE was used. Results in Fig. 3 show that JointDiff is most preferred, outperforming MoFlow (80%), U2Diff (65%), and w/o joint (53%). Removing ties increases the win rate over the ablated variant to 67%, indicating clear perceptual gains from joint modeling. JointDiff loses to GT in 44% of comparisons, with 24% ties, suggesting that many generated trajectories are difficult for users to distinguish from real ones.
@@ 5.2 Controllable Generation
Table 3: Consistency Analysis. Consistency between predicted events and trajectories, reported as maxmax / avgavg Acc $\uparrow$ over 20 samples. We compare our multinomial with the absorbing state framework.
Method Multinomial (Ours) Absorbing
NFL Bundesliga NBA NFL Bundesliga NBA
Ours w/o $\mathcal{G}$} .98 / .86 .97 / .80 .99 / .92 .97 / .80 .94 / .70 .99 / .89
Ours w ${\mathcal{G}}_{\text{WPG}}$ .96 / .84 .95 / .80 .99 / .92 .93 / .78 .91 / .68 .99 / .90
Ours w ${\mathcal{G}}_{\text{text}}$ .97 / .86 .96 / .81 - .95 / .81 .95 / .76 -
5.3 Consistency Analysis
We evaluate the consistency of the joint generation $\hat{\bm{\mathrm{X}}}_{0}=(\hat{\bm{\mathrm{Y}}}_{0},\hat{\bm{\mathrm{E}}}_{0})$. Using the same setup as in Table 2, we generate 20 samples per scene and compute the Acc metric between the predicted discrete possessor $\hat{\bm{\mathrm{E}}}_{0}$ and the threshold-based heuristic possessor extracted from the predicted trajectories $\hat{\bm{\mathrm{Y}}}_{0}$. For reference, we compare our multinomial diffusion with the commonly used absorbing state formulation (Levi et al., 2023; Li et al., 2025). For the absorbing, we use the publicly available implementation from Li et al. (2025) and set the absorption rate to $\lambda=0.01$ to match the magnitude of the discrete loss with the continuous one. Unlike the multinomial model, which enables refinement through all states at each denoising step, the absorbing mechanism freezes tokens once unmasked, preventing later correction (von Rütte et al., 2025). As shown in Table 3, our method achieves high consistency, particularly for the best sample (maxmax), and strong consistency on average (avgavg). The larger variance in NFL and Bundesliga aligns with their smaller training sets. Overall, this ablation shows that multinomial diffusion yields substantially more consistent predictions than the absorbing state approach in our dynamic domain. To assess the quality of the generations using absorbing state with respect to the ground truth, results for future generation are shown in Appendix D.3.5.
@@ 6 Conclusions
Acknowledgment. This work has been supported by the project GRAVATAR PID2023-151184OB-I00 funded by MCIU/AEI/10.13039/501100011033 and by ERDF, UE and by the Gov. of Catalonia under 2023 DI 00058.
@@ Appendix B Architecture
Figure 5: Model Architecture (extended version of Fig. 2). The light gray box represents the JointDiff architecture, and STB refers to the Social-Temporal Block.
@@
  • Temporal Mamba (TM): As described in Capellera et al. (2025), TM applies bidirectional Mamba modules (Gu and Dao, 2023) independently to each agent and sums the results.
@@ C.1 Possessor threshold
Our methodology is based on the key observation that a ball’s trajectory is primarily linear when it is not in any player’s possession (e.g., during a pass or a shot). Conversely, a player taking possession or influencing the ball’s path will induce a significant change in its direction. To quantify this, we analyzed the change in the ball’s direction by computing the angle between consecutive ball velocity vectors across our training datasets. For a range of distance thresholds from 0 to 3 meters, we calculated the average angle of change for the ball only during periods when it was outside of that specific threshold distance from all players. The results are shown in Fig. 6.
@@ C.2.1 NFL Dataset
  • Data Source: We process raw tracking data from the NFL Big Data Bowl, adopting the data splits and experimental settings from Xu and Fu (2025).
@@ D.1 Baselines
  • Sports-Traj (Xu and Fu, 2025): We used the official repository with the same hyperparameters as in their benchmark. We observed, consistent with the authors’ checkpoints in another benchmark, that the 20 sampled modes were nearly identical.
@@ D.2.1 Human-based Metrics
For each of the first 128 past observations in our test set, we formed a random pair from these selected samples. An example of the interface is shown in Fig. 7.
@@ D.2.3 Completion generation
We benchmark our approach on general trajectory completion using the datasets and experimental setup from Xu and Fu (2025), which includes the Basketball-U, Football-U, and Soccer-U datasets. In this task, a pre-defined mask using different strategies is applied to a scene, and the model must complete the missing observations. We note that while our NFL dataset uses the same data splits as Football-U, it employs a different masking strategy. For comparison, we reuse the evaluation table from Xu and Fu (2025), also reported in Capellera et al. (2025), which includes a wide range of baselines. These baselines span statistical methods: Mean, Median, Linear Fit; vanilla models: LSTM (Hochreiter and Schmidhuber, 1997), Transformer (Vaswani et al., 2017)); and advanced methods: MAT (Zhan et al., 2019), Naomi (Liu et al., 2019), INAM (Qi et al., 2020), SSSD (Alcaraz and Strodthoff, 2022), GC-VRNN (Xu et al., 2023), Sports-Traj (Xu and Fu, 2024), and U2Diff (Capellera et al., 2025).
Results are reported in Table 5. We use the ADE metric as defined in Xu and Fu (2025), which we rename as BADE to reflect its dependence on batch size and distinguish it from the standard individual-level ADE or the scene-level SADE. The BADE metric is defined as:
@@
where $y^{b}_{t,n}$ is the ground truth 2D spatial location of agent $n$ at timestep $t$ in scene $b$, $\hat{y}^{b}_{t,n}$ is its estimation, and $m_{t,n}^{b}$ is a value from the binary mask $\mathbf{M}$ where a value of 0 indicates a location to be predicted. In their setting, Xu and Fu (2025) use $B=128$. We also report the standard SADE metric. The table presents the minimum (minmin) values across 20 generated modes. While the minimum mode for BADE is selected across the entire batch of scenes, the minimum mode for SADE is selected for each scene independently, making it independent of batch size and mode ordering.
@@
Table 5: Completion Generation. We report the minmin for BADE (ADE in (Xu and Fu, 2025)) and SADE over 20 generated modes.
@@
D.2.4 Future generation
We further evaluate our method on future trajectory prediction using the basketball dataset, referred to as NBA12/13, and protocol from Li et al. (2021). This setting enables comparison against temporal autoregressive baselines such as NRI (Kipf et al., 2018), dNRI (Graber and Schwing, 2020), and GRIN (Li et al., 2021); GAN-based models such as Social-GAN (Gupta et al., 2018); and a transformer-based approach, FQA (Kamra et al., 2020). Results are presented in Table 6, which reports metrics equivalent to our minmin SADE and SFDE over 100 samples. JointDiff achieves strong performance relative to these forecasting models while operating as a more general trajectory completion framework, benefiting in particular from the non-autoregressive nature of diffusion along the temporal dimension.
Table 6: Future Generation on NBA12/13. We report the minmin for SADE and SFDE (ADE and FDE in (Li et al., 2021), respectively) over 100 generated modes.
Method NBA12/13 (Feet)
SADE $\downarrow$ SFDE $\downarrow$
NRI 2.10 4.56
dNRI 2.02 4.52
FQA 2.42 4.81
Social-Gan 1.88 3.64
GRIN 1.72 3.59
JointDiff (Ours) 1.36 2.52
@@ D.3.2 Lambda in Joint Loss
D.3.3 Guidance Strength
As stated in the main paper, during inference we utilize a single forward pass corresponding to the conditioned output, which corresponds to using $w=0$ in CFG (Ho and Salimans, 2022). With our DDIM sampling, this choice yields optimal performance, while extreme values degrade results. This setting also avoids the extra cost of dual forward passes. In Table 9 we report the same SADE / SFDE metrics as in Table 2 with ${\mathcal{G}}_{\text{text}}$ for NFL and Bundesliga datasets. CFG training enables a unified model supporting both controllable ($w=0$) and non-controllable (w=1w=-1) generation without notable degradation compared to our dedicated non-controllable model (Table 2, Ours w/o $\mathcal{G}$}).
Table 9: Guidance Weight $w$. We report SADE and SFDE metrics (minmin / avgavg) over 20 generated modes.
$w$ NFL (yards) Bundesliga (meters)
SADE $\downarrow$ SFDE $\downarrow$ SADE $\downarrow$ SFDE $\downarrow$
-1.0 2.44 / 3.47 5.75 / 8.59 2.51 / 3.70 5.10 / 8.33
-0.5 2.21 / 3.11 5.09 / 7.58 2.12 / 2.84 4.19 / 6.06
0.0 2.19 / 3.09 5.04 / 7.52 2.08 / 2.72 4.09 / 5.68
0.5 2.19 / 3.10 5.04 / 7.56 2.12 / 2.73 4.18 / 5.66
1.0 2.21 / 3.15 5.10 / 7.68 2.21 / 2.81 4.40 / 5.84
2.0 2.31 / 3.31 5.34 / 8.07 2.49 / 3.07 5.05 / 6.41
3.0 2.46 / 3.53 5.69 / 8.58 2.80 / 3.38 5.71 / 7.04
D.3.4 T5 Encoder
We compared the performance using T5-Small, T5-Base, and T5-Large encoders and observed only minor performance differences across the board. As shown in Table 10, T5-Base offers the best overall balance, achieving the lowest minmin / avgavg SADE and SFDE metrics on the NFL dataset, and competitive results on Bundesliga.
T5-Large performs slightly worse than T5-Base, which is likely due to the stronger dimensionality compression required to map its $d=1024$ dimensional embeddings down to our model’s 256-dimensional hidden space. Overall, the encoder size has minimal impact on the model’s performance, but the results suggest that excessive compression of a very large embedding space can lead to a slight degradation in trajectory prediction accuracy.
Table 10: T5 Encoder Size. We report SADE and SFDE metrics (minmin / avgavg) for different T5 encoder sizes ($d$), measured over 20 generated modes.
T5 Encoder ($d$) NFL (yards) Bundesliga (meters)
SADE $\downarrow$ SFDE $\downarrow$ SADE $\downarrow$ SFDE $\downarrow$
Small (512) 2.22 / 3.12 5.12 / 7.59 2.08 / 2.77 4.06 / 5.82
Base (768) 2.19 / 3.09 5.04 / 7.52 2.08 / 2.72 4.09 / 5.68
Large (1024) 2.24 / 3.25 5.14 / 7.88 2.10 / 2.75 4.11 / 5.74
D.3.5 Multinomial vs Absorbing
In the main paper, we compare the consistency of our multinomial discrete diffusion model with the absorbing state formulation in Table 3. To provide a more comprehensive assessment of generation quality, we extend this analysis using the same metrics as in Table 2 for controllable future generation. The results, presented in Table 11, compare our method when replacing the multinomial parameterization with an absorbingstate one (Ours w absorbing) against our original model. Across all tasks, the absorbing formulation exhibits consistently lower performance, which we attribute to reduced consistency between the generated trajectories and discrete events.
Table 11: Controllable Generation with Absorbing. This table compares our JointDiff model (using multinomial diffusion) against an ablation using absorbing state diffusion for discrete events (Ours w absorbing). As in Table2, we report the same metrics for both non-controllable (w/o $\mathcal{G}$}) and controllable (w ${\mathcal{G}}_{\text{WPG}}$, w ${\mathcal{G}}_{\text{text}}$) future generation tasks.
Method NFL (yards) Bundesliga (meters) NBA (meters)
Ours w absorbing
  w/o $\mathcal{G}$} 2.50 / 3.67 5.78 / 8.84 .76 / .54 2.72 / 4.13 5.39 / 8.93 .64 / .37 1.45 / 2.10 2.62 / 4.07 .73 / .44
  w ${\mathcal{G}}_{\text{WPG}}$ 2.57 / 3.84 5.72 / 8.92 .83 / .65 2.36 / 3.44 4.43 / 7.04 .75 / .49 1.28 / 1.89 2.25 / 3.66 .87 / .67
  w ${\mathcal{G}}_{\text{text}}$ 2.41 / 3.48 5.46 / 8.29 .85 / .72 2.29 / 3.21 4.34 / 6.44 .79 / .57 - - -
Ours
D.3.6 Additional Analysis
We present the last ablation analysis which includes the relevance of two important components in our approach. The first one is the importance sampling from Nichol and Dhariwal (2021), and the second is the number of Social Transformers layers within residual denoising block. Table 12 shows the results of our approach without the importance sampling (“w/o IS”) and with only one Social Transformer layer (“w 1 $\times$ ST”). Notice that the importance sampling is crucial when the size of the dataset is small, like the NFL and the Bundesliga. Using two Social Transformers also improve the results across the three datasets.